OpenStudy (anonymous):

please help 5/6 ÷ (-2/3)^2

OpenStudy (littlenugget):

Hi!

OpenStudy (littlenugget):

\[\frac{ 5 }{ 6 }\div \left(\begin{matrix}-2 \\ 3\end{matrix}\right)^{2}\]

OpenStudy (littlenugget):

So first to make this easier we are going to get rid of the exponent, how do we do that?

OpenStudy (littlenugget):

What does an exponent do? It multiplies the number by itself right?

OpenStudy (littlenugget):

So let's do that: \[\frac{ -2 }{ 3}\times \frac{ -2 }{ 3 }= \frac{ 4 }{ 9 }\]

OpenStudy (anonymous):

thanks and i need the work very fast sorry :P

OpenStudy (littlenugget):

ok lol ;)

OpenStudy (littlenugget):

so now you have 5/6 divided by 4/9 and when you divide you just turn the second number upside down

OpenStudy (littlenugget):

so 5/6 times 9/4 = 45/24

OpenStudy (littlenugget):

simplify it by 3

OpenStudy (littlenugget):

15/8

OpenStudy (anonymous):

ok thanks :)

OpenStudy (littlenugget):

No prob, lol sorry if i took too long xD and welcome to openstudy!